# Optimizing Optimizers for Fast Gradient-Based Learning

### v1

@@ Abstract
We lay the theoretical foundation for automating optimizer design in gradient-based learning. Based on the greedy principle, we formulate the problem of designing optimizers as maximizing the instantaneous decrease in loss. By treating an optimizer as a function that translates loss gradient signals into parameter motions, the problem reduces to a family of convex optimization problems over the space of optimizers. Solving these problems under various constraints not only recovers a wide range of popular optimizers as closed-form solutions, but also produces the optimal hyperparameters of these optimizers with respect to the problems at hand. This enables a systematic approach to design optimizers and tune their hyperparameters according to the gradient statistics that are collected during the training process. Furthermore, this optimization of optimization can be performed dynamically during training.
@@ 1 Introduction
  • Optimizers and their hyperparameters can be dynamically tuned or even be replaced by better ones according to the intermediate probes from the gradients in the middle of training.
  • By reverse engineering commonly used optimizers, we draw the landscape of optimizers that have driven the success of machine learning (robbins1951stochastic; kingma2015adam; loshchilov2017decoupled; gupta2018shampoo; martens2015optimizing) into a single picture.
@@ 2 Optimal stateless optimizers
Consider the following setup: Given a data distribution $\pi$, training samples $x\sim\pi$ produce the gradients $g=\nabla_{\theta}{\mathcal{L}}(\theta,x)$. The gradient moment is defined as:
$\Sigma\;\coloneqq\;\mathbb{E}[g\,g^{\top}],$ (3)
where $\mathbb{E}$ denotes the expectation over the gradient distribution. Note that $\Sigma$ is a symmetric and positive semidefinite (PSD) matrix of shape $d\times d$. For any symmetric PSD optimizer $Q\in\mathbb{S}_{+}^{d}$ of shape $d\times d$, let us define its learning power as an expected quadratic form of the gradients:
@@
Solving this without any additional constraint ends up with arbitrarily large eigenvalues of $Q$, which corresponds to arbitrarily large learning rates. This is certainly infeasible in practice. Real problems give us several reasons that make this “ideal solution” unrealizable: finite precision of our machines, curvature of non-convex loss landscapes, stochastic nature of subset gradients, etc. All of them restrict the ability of gradient estimates $g$ to represent the global geometry of the loss landscape over the parameter space. Taking a large step in the parameter space beyond the regions where $g$ remains explainable leads to unexpected, and usually fatal, behaviors.
This calls for additional constraints on the optimizer $Q$. AdaReg (gupta2017unified) considers similar problem as P2 but with an indirect regularization term $\Phi(Q)$ on $Q$. Instead, we allow the engineer to choose the trust region ${\mathcal{Q}}\subseteq\mathbb{S}_{+}^{d}$ that circumscribes the feasible set directly, leading to a family of exact solutions. The following theorem makes this precise:
Theorem 1 (Optimal stateless optimizers under convex constraints).
@@
    $P^{\star}\;=\;\delta_{{\mathcal{Q}}}^{*}\;=\;\gamma_{{\mathcal{Q}}^{\circ}}\qquad\textnormal{and}\qquad\gamma_{{\mathcal{Q}}}^{*}\;=\;\delta_{{\mathcal{Q}}^{\circ}},\$(5)
@@
Theorem 1 states that setting the family of available trust regions $\mathcal{Q}$} determines the optimal optimizer $Q^{\star}$ and its associated hyperparameters through the solutions to the convex optimization problem P2. In particular, consider the following four types of trust regions:
@@
Instantiating the construction from Theorem 1 on each of these families, we obtain the closed-form optimal optimizer $Q^{\star}$ and the corresponding optimal learning power $P^{\star}$.
Corollary 2 (Closed-form solutions for common optimizer families).
@@
Again, the proof is in Appendix G. These analytic solutions reveal how different types of optimal optimizers $Q^{\star}$ emerge from the choice of trust region geometry $\mathcal{Q}$}. Specifically, we see that:
@@
These optimal optimizers are proportional to the gradient moment $Q^{\star}\propto\Sigma$. We can further project this general class into special geometries to obtain the optimal hyperparameters for various types of optimizers as in Corollary 5.
@@
The spectral trust region ${\mathcal{Q}}_{\textnormal{S}}(\tau,\lambda)$ results in a water-filling optimizer that concentrates the parameter update rates into the largest available principal component of the data moment $\Sigma$ up to a per-mode cap $\lambda$ (similar to gradient clipping), sequentially, until the total budget $\tau$ is reached (similar to learning rate scheduling). This shows how algorithmic tricks can be represented as trust regions.
@@
The optimal solutions of this family allocate uniform power across eigendirections of the gradient moment $\Sigma$, whitening the gradient statistics by projecting onto the support of $\Sigma$ with a constant scaling factor. When $\Sigma$ is a Fisher information matrix, this represents natural gradient descent (amari1998natural). Generally, this encompasses full-matrix AdaGrad (duchi2011adaptive; agarwal2019efficient), K-FAC (martens2015optimizing), and Shampoo (gupta2018shampoo).
@@
Corollary 2 shows that these coordinate-wise optimal optimizers set up their weights that scale with the coordinate-wise gradient variance $\sigma_{j}=\Sigma_{jj}$ and inversely with the pre-defined costs $c_{j}$, leading to optimal hyperparameters for various types of existing optimizers such as Adam optimizer (kingma2015adam) as in Corollary 6. Various types of well-used optimizers are categorized by the choice of costs $c_{j}$, e.g., $c_{j}=1$ gives simple gradient descent, $c_{j}=\text{EMA}(g_{j}^{2})^{1/2}$ gives diagonal AdaGrad (duchi2011adaptive) or RMSProp-style optimizer (tieleman2012lecture).
The behaviors of different types of optimizers are visualized in Figure 2. Figure 2a and 2d show that our analytically found optimal optimizer is the fastest among all hyperparameter settings under the same Frobenius family. This demonstrates that we can systematically choose the optimal hyperparameters according to the gradient statistics as the theory suggests. On the other hand, Figure 2b, 2c, 2e, and 2f highlight how optimizers from different families can outperform others in their specialized domains. This reminds us of the notorious no free lunch theorem in optimization (wolpert1997no): “no single algorithm is universally superior”. With our convex optimization framework which associates the best algorithm that matches a given task, we can now update this catchphrase into: “Under the engineer’s choice of trust region $\mathcal{Q}$}, the optimal optimizer $Q^{\star}$ is determined by the task’s gradient statistics $\Sigma$”.
3 Optimal dynamic optimizers with state variables
Up to this point, we have focused on the simple stateless optimizers, which are less used in real-world applications. In practice, optimizers have memory, often in the form of momentum, in order to stabilize the learning process from stochastic gradients and non-convex loss landscapes. We now extend our framework by letting the optimizer $Q[n]$ be a causal dynamical operator: a filter that translates gradient history $g[n]$ into instantaneous parameter velocity $\dot{\theta}[n]$ (lee2024grokfast), where $n\in\{0,1,2,\ldots\}$ representing (discrete) training steps. Define a dynamic optimizer as an LTI filter with a symmetric matrix impulse response $Q[n]\in\mathbb{R}^{d\times d}$. This operates as a causal convolution:
$\dot{\theta}[n]\;=\;(Q*g)[n]\;\coloneqq\;\sum_{k=0}^{\infty}Q[k]\,g[n-k].$ (6)
We use the Hilbert norm $\langle\cdot,\cdot\rangle_{{\mathcal{H}}}$ instead of the Frobenius norm $\langle\cdot,\cdot\rangle_{F}=\operatorname{Tr}(\cdot^{\top}\cdot)$ we used in Section 2.
$\|Q\|_{{\mathcal{H}}}^{2}\;\coloneqq\;\sum_{n=0}^{\infty}\operatorname{Tr}(Q[n]^{\top}Q[n])\;<\;\infty,\qquad\langle Q_{1},Q_{2}\rangle_{{\mathcal{H}}}\;\coloneqq\;\sum_{n=0}^{\infty}\operatorname{Tr}(Q_{1}[n]^{\top}Q_{2}[n]).$ (7)
In addition, for theoretical analysis, gradient processes $g[n]\in\mathbb{R}^{d}$ are assumed to be wide-sense stationary (WSS). The moments are then the autocorrelations:
@@ 3 Optimal dynamic optimizers with state variables
for $k\geq 0$. The instantaneous learning power is, again, an inner product (details in Appendix F.1):
$P(Q;n)\coloneqq\mathbb{E}\big[g[n]^{\top}\dot{\theta}[n]\big]=\mathbb{E}\!\left[g[n]^{\top}\sum_{k=0}^{\infty}Q[k]\,g[n-k]\right]=\sum_{k=0}^{\infty}\operatorname{Tr}\!\left(Q[k]^{\top}R[k]\right)=\langle Q,R\rangle_{{\mathcal{H}}}.$ (9)
@@
Theorem 3 (Optimal dynamic optimizers under convex constraints).
@@
The proof is similar to the stateless case, and is provided in Appendix G like all other proofs for this section. Theorem 3 formalizes how the optimal dynamic optimizer $Q^{\star}$ equalizes the learning power across different frequencies as a function of the convex trust region $\mathcal{Q}$}. All the closed-form solutions from Corollary 2 can also be directly lifted to the dynamic framework, as elaborated in Appendix F.3. Instead of redundantly repeating these closed-form solutions, we discuss how they are connected to well-used optimizers in practice. As we will see in Corollaries 5 and 6, solving P3 using Theorem 3 often produces general dynamic optimizers with infinite impulse responses (IIR) $Q[n]$, whose implementations require infinite memory for the optimizer states. In practice, we often restrict ourselves to simpler, realizable families of optimizers, e.g., those with EMA-based momenta. The following lemma justifies this post-projection of optimal optimizers from general trust regions to more restrictive geometries with fewer controllable hyperparameters.
Lemma 4 (Conservation of optimality under projection).
@@
${\mathcal{Q}}^{\star}(R)\coloneqq\arg\max_{Q\in{\mathcal{Q}}}\langle Q,R\rangle_{{\mathcal{H}}},\qquad{\mathcal{Q}}_{\mathcal{C}}^{\star}(R)\coloneqq\arg\max_{Q\in{\mathcal{Q}}\cap{\mathcal{C}}}\langle Q,R\rangle_{{\mathcal{H}}}.$ (10)
@@
The projection cone $\mathcal{C}$} represents a desired property of the set of optimizers. If $\mathcal{C}$} is in a nice shape, e.g., a set of optimizers with a momentum ${\mathcal{C}}={\mathcal{C}}_{\text{1p}}\coloneqq\{Q[n]=\eta(1-\beta)\beta^{n}I:\eta\geq 0,0<\beta<1\}$, then we can first solve the easier unconstrained problem for general $\mathcal{Q}$} as in Corollary 2 and then project the resulting solutions to the cone $\mathcal{C}$} to obtain the final optimal solution over the target constrained problem ${\mathcal{Q}}\cap{\mathcal{C}}$. This extends our framework to many practical optimizers. Now we are ready to find the optimal hyperparameters for real optimizers in use.
Corollary 5 (Instantaneously optimal SGD+Momentum).
${\mathcal{Q}}_{\textnormal{F}}(B)\;\coloneqq\;\{Q:\|Q\|_{\mathcal{H}}\leq\sqrt{B}\},\quad{\mathcal{C}}_{\textnormal{1p}}\;\coloneqq\;\{Q_{\eta,\beta}[n]=\eta(1-\beta)\beta^{n}I:\eta\geq 0,0<\beta<1\}.$ (11)
Given gradients $g[n]$, define $m_{\beta}[n]\coloneqq\sum_{k=0}^{\infty}\beta^{k}g[n-k]$ as the unnormalized momentum at time $n$ with momentum parameter $\beta$. Then the optimal solution of problem P3 under the trust region ${\mathcal{Q}}_{\textnormal{F}}(B)\cap{\mathcal{C}}_{\textnormal{1p}}$ is an SGD+Momentum optimizer with optimal hyperparameters:
$\beta^{\star}[n]\;=\;\arg\underset{\beta\in(0,1)}{\max}\sqrt{1-\beta^{2}}\,\,\mathbb{E}[g[n]^{\top}m_{\beta}[n]],\quad\eta^{\star}\;=\;\left(\frac{B\,(1+\beta^{\star})}{d\,(1-\beta^{\star})}\right)^{\!1/2},$ (12)
where $d$ is the dimension of the parameter space.
The learning rate $\eta^{\star}$ scales to saturate the budget $B$.
Corollary 6 does the same to Adam (kingma2015adam). For Adam, the existence of a time-varying divisor $\operatorname{EMA}(g^{2},\beta_{2})^{-1/2}$ slightly complicates the derivation by making the optimizer time-varying.
Corollary 6 (Instantaneously optimal Adam/AdamW).
${\mathcal{Q}}_{\textnormal{D}}(B,c)$ $\;\coloneqq\;\{\operatorname{diag}(q_{j}):\textstyle\sum_{j}c_{j}\textstyle\sum_{k\geq 0}|q_{j}[k]|^{2}\leq B\},$ (13)
${\mathcal{C}}_{\textnormal{1p}}(c)$ $\;\coloneqq\;\{Q_{\eta,\beta_{1}}[n]=\operatorname{diag}{(\eta(1-\beta_{1})\beta_{1}^{n}/c_{j})}:\eta\geq 0,0<\beta_{1}<1\}.$ (14)
$m_{\beta_{1},j}[n]\;=\;\beta_{1}m_{\beta_{1},j}[n-1]+(1-\beta_{1})g_{j}[n],\quad v_{\beta_{2},j}[n]\;=\;\beta_{2}v_{\beta_{2},j}[n-1]+(1-\beta_{2})g_{j}^{2}[n],$ (15)
where $\beta_{1},\beta_{2}\in(0,1)$ are hyperparameters. Define the cost $c_{j}\;=\;v_{\beta_{2},j}^{1/2}$. Then the optimal solution of problem P3 under ${\mathcal{Q}}_{\textnormal{D}}(B,c)\cap{\mathcal{C}}_{\textnormal{1p}}(c)$ is an Adam optimizer with optimal hyperparameters:
$(\beta_{1}^{\star}[n],\beta_{2}^{\star}[n])=\arg\underset{\beta_{1}\in(0,1),\beta_{2}\in(0,1)}{\max}a(\beta_{1},\beta_{2})\,\mathbb{E}[g[n]^{\top}u_{\beta_{1},\beta_{2}}[n]],\quad\eta^{\star}=\sqrt{B}\,a(\beta_{1}^{\star},\beta_{2}^{\star}),$ (16)
where $u_{\beta_{1},\beta_{2}}\coloneqq m_{\beta_{1}}/v_{\beta_{2}}^{1/2}$ is the Adam update and $a(\beta_{1},\beta_{2})\coloneqq\sqrt{(1+\beta_{1})/((1-\beta_{1})\textstyle\sum_{j}(1/c_{j}))}$ is the normalization factor.
Table 1: Demonstration of Corollary 5 for SGD with momentum. Best baseline at $\beta=0.9$. mean $\pm$ std.
@@
Table 2: Demonstration of Corollary 6 for Adam. Best baseline at $(\beta_{1},\beta_{2})=(0.8,0.999)$. mean $\pm$ std.
@@
Figure 3: Demonstration of Corollaries 5 and 6. Our instantiations of optimal optimizers are compared with baselines having fixed hyperparameters on the CIFAR-100 dataset (krizhevsky2009learning) with ResNet-18 (he2016deep), following the standard settings of he2016deep. The error bars indicate the mean and standard deviation over 10 runs. Our instantiation shows better performance than every baseline optimizer with fixed hyperparameters, without relying on heavy workload of manual hyperparameter tuning.
Corollaries 5 and 6 show that the optimal hyperparameters for both SGD+Momentum and Adam are the ones that maximize the expected weighted cosine similarity between the gradients and the optimizer responses. The optimal hyperparameters are achieved when the corresponding optimizer aligns best with the gradient distribution for each time step $n$.
Furthermore, by deriving the SGD and Adam/AdamW optimizers from our framework, we gain additional insights into these well-known optimizers. For example, Adam/AdamW can be interpreted as an optimal 1-pole approximation of the dynamic diagonal optimizer with cost $c=v^{1/2}$. Similarly, we can reverse engineer various other optimizers, including Gauss-Newton, natural gradient descent (amari1998natural), K-FAC (martens2015optimizing), Shampoo (gupta2018shampoo), and Muon (liu2025muon), into our framework, as shown in Table 7 and in Appendix E. This classifies practical optimizers, reveals their hidden design principles, and provides a systematic way to determine optimal hyperparameters for these optimizers.

### v2

# Let the Optimizers Optimize Themselves

@@ Abstract
We lay the theoretical foundation for automating optimizer design in gradient-based learning. Based on the greedy principle, we formulate the problem of designing optimizers and their hyperparameters as maximizing the instantaneous decrease in loss. By treating an optimizer as a function that translates loss gradient signals into parameter motions, the problem reduces to a family of convex optimization problems over the space of optimizers. Solving these problems under various constraints not only recovers a wide range of popular optimizers as closed-form solutions, but also produces the optimal hyperparameters of these optimizers with respect to the problems at hand. This enables a systematic approach to design optimizers and tune their hyperparameters dynamically according to the gradient statistics that are collected during the training process. Furthermore, this optimization of optimization can be performed dynamically during training.
No Free Lunch, Optimization, Gradient-Based Learning, Convex Optimization, Dynamic Optimizer, First-Order Method, Power Allocation, Quadratic Form, AutoML, Machine Learning, ICML
@@ 1 Introduction
  • Optimizers and their hyperparameters can be dynamically updated according to the intermediate probes from the gradients in the middle of training.
  • By reverse engineering commonly used optimizers, we draw the landscape of optimizers that have driven the success of machine learning, from SGD (Robbins and Monro, 1951) and Adam (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) to Muon (Liu and others, 2025) into a single picture, as fully summarized in Table 7.
@@ 2 Optimal Stateless Optimizers
Given a data distribution $\pi$, training samples $x\sim\pi$ produce the gradients $g=\nabla_{\theta}{\mathcal{L}}(\theta,x)$. The gradient moment is a symmetric and positive semidefinite (PSD) matrix of shape $d\times d$ defined as: $\Sigma\coloneqq\mathbb{E}[gg^{\top}]$, where $\mathbb{E}$ denotes the expectation over the gradient distribution. For any symmetric PSD optimizer $Q\in\mathbb{S}_{+}^{d}$ of shape $d\times d$, let us define its learning power as an expected quadratic form of the gradients:
@@
Solving this without any additional constraint ends up with arbitrarily large eigenvalues of $Q$, corresponding to arbitrarily large learning rates. This is certainly infeasible in practice. Real problems give us several reasons that make this “ideal solution” unrealizable: finite precision of our machines, curvature of non-convex loss landscapes, stochastic nature of subset gradients, etc. All of them restrict the ability of gradient estimates $g$ to represent the global geometry of the loss landscape over the parameter space. Taking a large step in the parameter space beyond the regions where $g$ remains explainable leads to unexpected, and usually fatal, behaviors. This calls for additional constraints on the optimizer $Q$. AdaReg (Gupta et al., 2017) considers similar problem as P2 but with an indirect regularization term $\Phi(Q)$ on $Q$. Instead, we allow the engineer to choose the trust region ${\mathcal{Q}}\subseteq\mathbb{S}_{+}^{d}$ that circumscribes the feasible set directly, leading to a family of exact solutions. The following theorem makes this precise:
Theorem 2.1 (Optimal stateless optimizers under convex constraints).
@@
    $P^{\star}\;=\;\delta_{{\mathcal{Q}}}^{*}\;=\;\gamma_{{\mathcal{Q}}^{\circ}}\quad\textnormal{and}\quad\gamma_{{\mathcal{Q}}}^{*}\;=\;\delta_{{\mathcal{Q}}^{\circ}},\$(4)
@@
Theorem 2.1 states that setting the family of available trust regions $\mathcal{Q}$} determines the optimal optimizer $Q^{\star}$ and its associated hyperparameters through the solutions to the convex optimization problem P2. In particular, consider the following four types of trust regions:
@@
Instantiating the construction from Theorem 2.1 on each of these families, we obtain the closed-form optimal optimizer $Q^{\star}$ and the corresponding optimal learning power $P^{\star}$.
Corollary 2.2 (Closed-form solutions for common optimizer families).
@@
Again, the proof is in Appendix G. These analytic solutions reveal how different types of optimal optimizers $Q^{\star}$ emerge from the choice of trust region geometry $\mathcal{Q}$}. Specifically,
@@
These optimal optimizers are proportional to the gradient moment $Q^{\star}\propto\Sigma$. We can further project this general class into special geometries to obtain the optimal hyperparameters for various types of optimizers as in Corollary 3.3.
@@
The spectral trust region ${\mathcal{Q}}_{\textnormal{S}}(\tau,\lambda)$ results in a water-filling optimizer that concentrates the parameter update rates into the largest available principal component of the data moment $\Sigma$ up to a per-mode cap $\lambda$ (similar to gradient clipping), sequentially, until the total budget $\tau$ is reached (similar to LR scheduling). This shows how trust regions can represent algorithmic tricks.
@@
The optimal solutions of this family allocate uniform power across eigendirections of the gradient moment $\Sigma$, whitening the gradient statistics by projecting onto the support of $\Sigma$ with a constant scaling factor. When $\Sigma$ is a Fisher information matrix, this represents natural gradient descent (Amari, 1998). Generally, this encompasses full-matrix AdaGrad (Duchi et al., 2011; Agarwal et al., 2019), K-FAC (Martens and Grosse, 2015), and Shampoo (Gupta et al., 2018).
@@
Corollary 2.2 shows that these coordinate-wise optimal optimizers set up their weights that scale with the coordinate-wise gradient variance $\sigma_{j}=\Sigma_{jj}$ and inversely with the pre-defined costs $c_{j}$, leading to optimal hyperparameters for various types of existing optimizers such as Adam optimizer (Kingma and Ba, 2015) as in Corollary 3.4. Various types of well-used optimizers are categorized by the choice of costs $c_{j}$, e.g., $c_{j}=1$ gives simple gradient descent, $c_{j}=\text{EMA}(g_{j}^{2})^{1/2}$ gives diagonal AdaGrad (Duchi et al., 2011) or RMSProp (Tieleman and Hinton, 2012).
The behaviors of different types of optimizers are visualized in Figure 2. Figure 2a and 2d show that our analytically found optimal optimizer is the fastest among all hyperparameter settings under the same Frobenius family. This demonstrates that we can systematically choose the optimal hyperparameters according to the gradient statistics as the theory suggests. On the other hand, Figure 2b, 2c, 2e, and 2f highlight how optimizers from different families can outperform others in their specialized domains. This reminds us of the notorious no free lunch theorem in optimization (Wolpert and Macready, 1997): “no single algorithm is universally superior”. With our convex optimization framework which associates the best algorithm that matches a given task, we can now update this catchphrase into: “Under the engineer’s choice of trust region $\mathcal{Q}$}, the optimal optimizer $Q^{\star}$ is determined by the task’s gradient statistics $\Sigma$”.
3 Optimal Dynamic Optimizers with States
Up to this point, we have focused on the simple stateless optimizers, which are less used in real-world applications. In practice, optimizers have memory, often in the form of momentum, in order to stabilize the learning process from stochastic gradients and non-convex loss landscapes. We now extend our framework by letting the optimizer $Q[n]$ be a causal dynamical operator: a filter that translates gradient history $g[n]$ into instantaneous parameter velocity $\dot{\theta}[n]$ (Lee et al., 2024), where $n\in\{0,1,2,\ldots\}$ representing (discrete) training steps. Define a dynamic optimizer as an LTI filter with a symmetric matrix impulse response $Q[n]\in\mathbb{R}^{d\times d}$. This operates as a causal convolution:
$\dot{\theta}[n]\;=\;(Q*g)[n]\;\coloneqq\;\sum_{k\geq 0}Q[k]\,g[n-k].$ (5)
We use the Hilbert norm $\langle\cdot,\cdot\rangle_{{\mathcal{H}}}$ instead of the Frobenius norm $\langle\cdot,\cdot\rangle_{F}=\operatorname{Tr}(\cdot^{\top}\cdot)$ we used in Section 2:
$\langle Q_{1},Q_{2}\rangle_{{\mathcal{H}}}\coloneqq\sum_{n\geq 0}\operatorname{Tr}(Q_{1}[n]^{\top}Q_{2}[n]),$ (6)
and let $\|Q\|_{{\mathcal{H}}}^{2}=\sum_{n\geq 0}\operatorname{Tr}(Q[n]^{\top}Q[n])<\infty$. In addition, for theoretical analysis, gradient processes $g[n]\in\mathbb{R}^{d}$ are assumed to be wide-sense stationary (WSS). The moments are then the autocorrelations:
@@ 3 Optimal Dynamic Optimizers with States
for $k\geq 0$. The instantaneous learning power is, again, an inner product (detailed derivation is in Appendix F.1):
$P(Q;n)\coloneqq\mathbb{E}\big[g[n]^{\top}\dot{\theta}[n]\big]=\langle Q,R\rangle_{{\mathcal{H}}}.$ (8)
@@
Theorem 3.1 (Optimal dynamic optimizers under convex constraints).
@@
The proof is similar to the stateless case, and is provided in Appendix G like all other proofs for this section. Theorem 3.1 formalizes how the optimal dynamic optimizer $Q^{\star}$ equalizes the learning power across different frequencies as a function of the convex trust region $\mathcal{Q}$}. All the closed-form solutions from Corollary 2.2 can also be directly lifted to the dynamic framework, as elaborated in Appendix F.3. Instead of redundantly repeating these closed-form solutions, we discuss how they are connected to well-used optimizers in practice. As we will see in Corollaries 3.3 and 3.4, solving P3 using Theorem 3.1 often produces general dynamic optimizers with infinite impulse responses (IIR) $Q[n]$, whose implementations require infinite memory for the optimizer states. In practice, we often restrict ourselves to simpler, realizable families of optimizers, e.g., those with EMA-based momenta. The following lemma justifies this post-projection of optimal optimizers from general trust regions to more restrictive geometries with fewer controllable hyperparameters.
Lemma 3.2 (Conservation of optimality under projection).
@@
${\mathcal{Q}}^{\star}(R)$ $\coloneqq\arg\max_{Q\in{\mathcal{Q}}}\langle Q,R\rangle_{{\mathcal{H}}},$ (9)
${\mathcal{Q}}_{\mathcal{C}}^{\star}(R)$ $\coloneqq\arg\max_{Q\in{\mathcal{Q}}\cap{\mathcal{C}}}\langle Q,R\rangle_{{\mathcal{H}}}.$ (10)
@@
The projection cone $\mathcal{C}$} represents a desired property of the set of optimizers. If $\mathcal{C}$} is in a nice shape, e.g., a set of optimizers with a momentum ${\mathcal{C}}={\mathcal{C}}_{\text{1p}}\coloneqq\{Q[n]=\eta(1-\beta)\beta^{n}I:\eta\geq 0,0<\beta<1\}$, then we can first solve the easier unconstrained problem for general $\mathcal{Q}$} as in Corollary 2.2 and then project the resulting solutions to the cone $\mathcal{C}$} to obtain the final optimal solution over the target constrained problem ${\mathcal{Q}}\cap{\mathcal{C}}$. This extends our framework to many practical optimizers. Now we are ready to find the optimal hyperparameters for real optimizers in use.
Table 1: Demonstration of Corollary 3.3 for SGD with momentum. Best baseline at $\beta=0.9$. mean $\pm$ std.
@@
Table 2: Demonstration of Corollary 3.4 for Adam. Best baseline at $(\beta_{1},\beta_{2})=(0.8,0.999)$. mean $\pm$ std.
@@
Figure 3: Demonstration of Corollaries 3.3 and 3.4. Our instantiations of optimal optimizers are compared with baselines having fixed hyperparameters on the CIFAR-100 dataset (Krizhevsky, 2009) with ResNet-18 (He et al., 2016), following the standard settings of (He et al., 2016). The error bars indicate the mean and standard deviation over 10 runs. Our instantiation shows better performance than every baseline optimizer with fixed hyperparameters, without relying on heavy workload of manual hyperparameter tuning.
Corollary 3.3 (Instantaneously optimal SGD+Momentum).
${\mathcal{Q}}_{\textnormal{F}}$ $\coloneqq\{Q:\|Q\|_{\mathcal{H}}\leq\sqrt{B}\},$ (11)
${\mathcal{C}}_{\textnormal{1p}}$ $\coloneqq\{Q[n]=\eta(1-\beta)\beta^{n}I:\eta\geq 0,0<\beta<1\}.$ (12)
Given gradients $g[n]$, define $m_{\beta}[n]\coloneqq\sum_{k=0}^{\infty}\beta^{k}g[n-k]$ as the unnormalized momentum at time $n$ with momentum parameter $\beta$. Then the optimal solution of problem P3 under the trust region ${\mathcal{Q}}_{\textnormal{F}}(B)\cap{\mathcal{C}}_{\textnormal{1p}}$ is an SGD optimizer with optimal hyperparameters: $\eta^{\star}=\left(\frac{B\,(1+\beta^{\star})}{d\,(1-\beta^{\star})}\right)^{\!1/2}$ and
$\beta^{\star}[n]\;=\;\arg\underset{\beta\in(0,1)}{\max}\sqrt{1-\beta^{2}}\,\,\mathbb{E}[g[n]^{\top}m_{\beta}[n]],$ (13)
where $d$ is the dimension of the parameter space.
The learning rate $\eta^{\star}$ scales to saturate the budget $B$.
Corollary 3.4 does the same to Adam (Kingma and Ba, 2015). For Adam, the existence of a time-varying divisor $\operatorname{EMA}(g^{2},\beta_{2})^{-1/2}$ slightly complicates the derivation by making the optimizer time-varying.
Corollary 3.4 (Instantaneously optimal Adam/AdamW).
${\mathcal{Q}}_{\textnormal{D}}$ $\coloneqq\{\operatorname{diag}(q_{j}):\textstyle\sum_{j}c_{j}\textstyle\sum_{k\geq 0}|q_{j}[k]|^{2}\leq B\},$ (14)
${\mathcal{C}}_{\textnormal{1p}}(c)$ $\coloneqq\big\{Q[n]=\operatorname{diag}\big(\tfrac{\eta(1-\beta_{1})\beta_{1}^{n}}{c_{j}}\big):\eta\geq 0,0<\beta_{1}<1\big\}.$ (15)
$m_{\beta_{1},j}[n]$ $=\beta_{1}m_{\beta_{1},j}[n-1]+(1-\beta_{1})g_{j}[n],$ (16)
$v_{\beta_{2},j}[n]$ $=\beta_{2}v_{\beta_{2},j}[n-1]+(1-\beta_{2})g_{j}^{2}[n],$ (17)
where $\beta_{1},\beta_{2}\in(0,1)$ are hyperparameters. Define the cost $c_{j}\;=\;v_{\beta_{2},j}^{1/2}$. Then the optimal solution of problem P3 under ${\mathcal{Q}}_{\textnormal{D}}(B,c)\cap{\mathcal{C}}_{\textnormal{1p}}(c)$ is an Adam optimizer with optimal hyperparameters: $\eta^{\star}=\sqrt{B}\,a(\beta_{1}^{\star},\beta_{2}^{\star})$ and
$(\beta_{1}^{\star},\beta_{2}^{\star})[n]=\arg\underset{\beta_{1},\beta_{2}\in(0,1)}{\max}a(\beta_{1},\beta_{2})\,\mathbb{E}[g[n]^{\top}u_{\beta_{1},\beta_{2}}[n]].$ (18)
where $u_{\beta_{1},\beta_{2}}\coloneqq m_{\beta_{1}}/v_{\beta_{2}}^{1/2}$ is the Adam update and $a(\beta_{1},\beta_{2})\coloneqq\sqrt{\tfrac{1+\beta_{1}}{(1-\beta_{1})\sum_{j}c_{j}^{-1}}}$ is the normalization factor.
Corollaries 3.3 and 3.4 show that the optimal hyperparameters for both SGD+Momentum and Adam are the ones that maximize the expected weighted cosine similarity between the gradients and the optimizer responses. The optimal hyperparameters are achieved when the corresponding optimizer aligns best with the gradient distribution for each time step $n$.
Furthermore, by deriving the SGD and Adam/AdamW optimizers from our framework, we gain additional insights into these well-known optimizers. For example, Adam/AdamW can be interpreted as an optimal 1-pole approximation of the dynamic diagonal optimizer with cost $c=v^{1/2}$. Similarly, we can reverse engineer various other optimizers, including Gauss-Newton, natural gradient descent (Amari, 1998), K-FAC (Martens and Grosse, 2015), Shampoo (Gupta et al., 2018), and Muon (Liu and others, 2025), into our framework, as shown in Table 7 and in Appendix E. This classifies practical optimizers, reveals their hidden design principles, and provides a systematic way to determine optimal hyperparameters for these optimizers.